OpenStudy (anonymous):

A family is building a pool that must be a rectangular prism and hold 1400ft³ and-3200ft³ of water. It must also be between 6-10ft deep. The family is also building a pool for their dog that has dimensions 1/4 the size of the family pool. I need to know what the dimensions for each pool should be...I can't come up with the right dimensions/water volume that works for both pools.

OpenStudy (anonymous):

must hold BETWEEN 1400 and 3200ft³ of water

OpenStudy (nottim):

is this calculus? OR advanced functions? Or functions?

OpenStudy (nottim):

helllooo

OpenStudy (nottim):

@i_hate_math_

OpenStudy (anonymous):

Geometry

OpenStudy (anonymous):

you have,\[1400\le lwh \le 3200\] and, for the dog pool, with each dimension 1/4 of the pool above,\[(1/4)^3lwh\]

OpenStudy (anonymous):

I still don't get it =/ I'm so confused

OpenStudy (anonymous):

but, we have a restriction on the depth, h:\[6\le h \le 10\] There isn't going to be a unique answer here. But choosing l=10 w=14 and h=10 works. The dog pool dimensions will be l=10/4, w=14/4 and h=10/4. There are many other possibilities which fit your restrictions.

OpenStudy (anonymous):

Ok..what I'm not understanding though, is how the dog pool is 1/4 the size of the family pool. Because shouldn't the volume be 1/4 as well as the dimensions?

OpenStudy (anonymous):

the dimensions of the dog pool are 1/4 the size of the regular pool. The volume would not be 1/4 of the regular pool. It would by (1/4)^3=1/64 of the regular pool.